1. 15% of students have a balance on their credit card greater than $ 5,500. Suppose you select 8 students to interview them regarding the use of your credit card (Binomial)

A. What is the probability that 5 of the students have a balance on their credit card exceeding $ 5,500?
b. That at least 3 have a balance on your credit card exceeding $ 5,500?
Answer #1

Solution:-

p=0.15, q=(1-p)=0.85, n=8

A.)

B.)

=1-0.8948

=0.1052